OpenStudy (anonymous):

Translate into an equation and show steps to solve. Marisa and Lisa earned a total of $65 babysitting during the month of November. Marisa earned $5 more than 1/2 of what Lisa earned. How much did they each earn.

OpenStudy (anonymous):

Let x = 1/2 of what Lisa earned. x + 5 + 2x = 65 3x = 65 - 5 3x = 60 x = 60/3 x = 20 Therefore, if Marisa earned x + 5, she earned 20 + 5 which is $25. Similarly, if Lisa earned 2x, she earned 2(20) which is $40.

OpenStudy (anonymous):

Thanks for the help.